Suppose a consumer consuming two commodities X and Y has the following utility function U= 10X0.4Y0.6. If price of good X and Y are 2 and 3 respectively and income constraint is Birr 50, then find

Find quantities of X and Y which maximize utility

Show how the rise in income to Birr 100 will affect the quantities of X and Y



1
Expert's answer
2021-08-06T05:10:56-0400

Solution:

To maximize utility: MRSxy ="\frac{MUx}{MUy} = \frac{Px}{Py}"

Derive MUx:

U = 10X0.4Y0.6


MUx = "\frac{\partial U} {\partial X}" = 4X-0.6Y0.4


Derive MUy:


MUy = "\frac{\partial U} {\partial Y}" = 6X0.4Y-0.4


MRSxy = "\frac{MUx}{MUy}" = "\frac{4X^{-0.6}Y^{0.4} }{6X^{0.4}Y^{-0.4} }"


MRSxy = "\frac{Px}{Py}"


"\frac{2Y^{0.8} }{3X }" = "\frac{2 }{3 }"


Y = X1.25

Budget constraint: I = PxX + PyY

50 = 2X + 3Y

Substitute the value of X in budget constraint to derive Y:

50 = 2X + 3(X1.25)

50 = 2X + 3X1.25

X = 7

Y = X1.25 = 71.25 = 12

Y = 12

The quantities of X and Y which maximize utility (Uxy) = (7, 12)

 

New budget constraint: 100 = 2X + 3Y

Y = X1.25

Substitute the value of X in budget constraint to derive Y:

100 = 2X + 3(X1.25)

100 = 2X + 3X1.25

X = 13

Y = X1.25 = 131.25 = 25

Y = 25

The new quantities of X and Y which maximize utility (Uxy) = (13, 25)


The rise in income to Birr 100, will enable the consumer to double the quantities of X and Y since the consumers will be having more funds to spare.